# Dynamical Dark Energy from Lattice Quantum Gravity

### v1

@@ Abstract
We study the behavior of the vacuum in Euclidean dynamical triangulations (EDT). Algorithmic improvements and better lattice spacing determinations allow us to test the properties of the emergent de Sitter geometries of our simulations to higher precision than previously possible. Although the agreement with de Sitter is good, the improved precision reveals deviations that can be interpreted as non-trivial vacuum dynamics, well-described by a cosmological constant that runs with scale. The simulations show that the dominant running is quadratic and that the scale can be identified with the Hubble rate. Several key cross-checks support this picture, including consistent results across multiple lattice spacings and the fact that covariant energy conservation is maintained. The parameters of the running are fully determined by simulations, enabling predictions when extrapolated to the scales relevant for our universe. This leads to a model for dark energy that is compatible with current observations, but which predicts deviations from $\Lambda$CDM at the ${\cal O}(10^{-3})$ level in cosmological observables that could be tested with future improvements in precision measurements.
@@ I Introduction
The machinery introduced here and in Ref. [19] thus enables a strong test of our simulation results against expectations from the semiclassical de Sitter solution. This new precision allows us to show that despite the good overall agreement with this picture, there are deviations from semiclassical de Sitter that can be associated with nontrivial vacuum dynamics. This dynamics is well-described by a power-law running of the cosmological constant, where the renormalization scale is associated with the Hubble scale. This coincides with one of the simplest of a family of running vacuum models studied by Solà and collaborators [24, 25, 26, 27, 28]. These models have been compared to observations, thus constraining the free parameters of the models and testing whether the new physics implied by these models can resolve the emerging tensions in observational cosmology [27, 29]. We are able to turn this around by fitting lattice data directly to a specific running vacuum model, thus determining the free parameters of the model from first principles via our calculations. We find that only one parameter, a dimensionless coupling, is compatible with a nonzero value. The evidence for this parameter being different from zero is compelling, with high statistical significance and good agreement across four lattice spacings, making it unlikely that this effect is merely a lattice artifact.
Remarkably, the model selected by our lattice calculation is compatible with covariant energy conservation, which is itself a consequence of diffeomorphism invariance. The sign of the effect seen here also ensures that it does not violate the null-energy condition. There are strong constraints on theories that are inconsistent with either of these conditions [30, 31, 32], so the fact that our lattice result does not run afoul of them is a significant point in its favor. Although our results are found on lattices that are no larger than several Planck volumes, it is interesting to ask what would happen if the model singled out by this work was assumed to hold throughout cosmic history. One consequence is that the running of the vacuum energy would lead to a large vacuum energy density in the early universe compared to the current era. It turns out that the computed size of the coupling of the model ensures that the effects of the vacuum on the evolution of the universe remain subdominant throughout cosmic history, thus making it compatible with present observational bounds. Nonetheless, the corrections to standard cosmology would not be negligible, with the effects of the running vacuum giving rise to deviations from the $\Lambda$CDM model at the part-per-mil level across many cosmological observables, including the dark energy equation of state.
@@ II.1 Euclidean Dynamical Triangulations
$Z_{E}=\sum_{T}\frac{1}{C_{T}}\left[\prod_{j=1}^{N_{2}}{\cal O}(t_{j})^{\beta}\right]e^{-S_{ER}}$ (3)
where $C_{T}$ is a symmetry factor that divides out the number of equivalent ways of labeling the vertices in the triangulation $T$. The term in brackets in Eq. (19) is a nonuniform measure term [20, 33], where the product is over all triangles, and ${\cal O}(t_{j})$ is the order of triangle $j$, i.e. the number of four-simplices to which the triangle belongs. This corresponds in the continuum to a nonuniform weighting of the measure in Eq. (1) by $\prod_{x}\sqrt{g}^{\beta}$, where we leave $\beta$ an adjustable parameter in the simulations.
@@
$S_{E}=-\frac{\sqrt{3}}{2}\pi\kappa N_{2}+N_{4}\left(\kappa\frac{5\sqrt{3}}{2}\mbox{arccos}\frac{1}{4}+\frac{\sqrt{5}}{96}\lambda\right),$ (6)
where $N_{i}$ is the number of simplices of dimension $i$. We rewrite the Einstein-Regge action in the simple form
@@ II.3 The phase diagram and recovering physical results
The tuning of the bare lattice parameters to the phase transition is performed as follows: The emergent shape of the lattice geometries differs markedly between the two phases, allowing us to use a measurement of the shape to identify the transition. To do this, we introduce the shelling function $N_{4}^{\rm shell}(\tau)$, which is defined as the number of four-simplices lying exactly a geodesic distance $\tau$ away from a source four-simplex, comprising a spherical shell with the source at its center. This function is averaged over multiple random sources on a given configuration and over all configurations of an ensemble. We account for autocorrelation errors present in the data for the shelling function by blocking the data before averaging. We study the variation of the error with block size, increasing the block size until the standard error obtained from a jackknife average over configurations no longer increases. The height of the peak of the shelling function $N_{4,\mathrm{peak}}^{\rm shell}$ varies significantly across the phase transition and thus serves as a good order parameter for studying the phase diagram.
@@
As described in Ref. [19], there is a “knee” in the plots at each volume, just before the slope becomes large and negative, that marks the onset of the phase transition. We take a fixed value of $\kappa_{2}$ to define our nominal lattice spacing, and we then tune $\beta$ by matching the rescaled peak heights in the region just to the left of the phase transition. The $\beta$ value is chosen so that the tuned point is sufficiently far from the phase transition that there is no evidence of the tunneling between metastable states that is expected in the vicinity of a first-order phase transition. Adjusting the rescaling of the peak height so that the location of the “knee" is at the same height across different volumes leads to a $D_{H}$ close to the physical value of four. We thus assume that the peak height rescaling is four-dimensional, and we choose $\beta$ as a function of volume so that the rescaled peak heights match the rescaled peak height of one of our smaller volumes near the transition. The choice of that smaller volume varies across the different lattice spacings, with its corresponding peak height given as the cyan band in each plot of Fig. 2. It is the 8k volume at $\kappa_{2}=2.45$, the 8k volume at $\kappa_{2}=3.0$, the 8k volume at $\kappa_{2}=3.4$, and the 32k volume at $\kappa_{2}=3.8$. As we show in Section III, this collection of ensembles passes a broad set of consistency checks, including the recovery of the semiclassical limit, suggesting that this tuning procedure is a valid one.
Figure 2: Rescaled peak heights for multiple volumes at four lattice spacings. The horizontal cyan line represents the peak height to which the tuning is performed for all volumes at that lattice spacing.
@@ III.1 The classical de Sitter solution
where $\ell$ is the dual lattice spacing and $\ell_{V}$ is the unit length derived from taking the cube root of the volume factor as measured by the shelling function. These units of measurement can be related via a comparison of the total four-volume of a thin slice of de Sitter space measured either by summing the volumes of four-simplices making up that slice, or by taking the integral over a four-dimensional shell. This leads to the relation
@@ III.3.1 $a_{\rm lat}/\ell$ from Method 1
Figure 4: The rescaled shelling function $n_{4}(\rho)$ at different lattice spacings. The black curve is the classical de Sitter solution. All curves have been rescaled to overlap in the region that matches the classical solution. The asymmetry at large Euclidean time decreases as the lattice spacing gets finer, so that the lattice results approach the classical curve.
@@ III.4 Determining Newton’s constant and setting the lattice spacing
Table 17, Table 18, Table 19, and Table 20 show the results for $|s_{G}|$ from all of our scaling studies. The first error is statistical, and the second is an estimate of the systematic error due to varying the $\kappa_{4}$ data points included in the fits. Adding these errors in quadrature and constructing $G/a_{\rm lat}^{2}$ from Eq. (28) leads to the values of $G/a_{\rm lat}^{2}$ shown in Fig. 11, plotted as a function of $1/N_{4}$. We have determined $G/a_{\rm lat}^{2}$ at multiple volumes at each lattice spacing, allowing us to use the infinite-volume limit of $G/a_{\rm lat}^{2}$, evaluated separately for each $\kappa_{2}$ value, to determine a relative lattice spacing. In previous work [21] there was only one lattice spacing available with multiple volumes, so this was not possible. In that case, the relative lattice spacing was determined using the diffusion process of a random walk. In revisiting the return probability to obtain the relative lattice spacing on our new ensembles, we find that there is significant ambiguity in rescaling the return probability curve, given how much the shape changes with lattice spacing. Using our $G/a_{\rm lat}^{2}$ for this purpose removes that ambiguity, while also being systematically improvable with more computing resources. We take the $\kappa_{2}=3.0$ ensembles to define our fiducial lattice spacing, so that $a_{\rm rel}$ at this lattice spacing is by definition one. The other lattice spacings are measured relative to this, being inversely proportional to the square root of the ratio of their $G/a_{\rm lat}^{2}$ values. The value of $G/a_{\rm lat}^{2}$ at $\kappa_{2}=3.0$ then sets our absolute lattice spacing, allowing us to convert our measurements to units of the Planck length.
@@ IV.1 Theory and motivation
Our lattice geometries are well-described by the semi-classical Euclidean de Sitter solution, which is solid evidence that EDT could define a non-perturbative formulation of quantum gravity. Interestingly, a more careful study shows that there are corrections to this picture. As pointed out in Section III.3.2, we find that the cosmological constant is not strictly constant as a function of Euclidean time, particularly on our smaller volume ensembles. Figure 12 shows the cosmological constant inferred from Eq. 36 as a function of Euclidean time $\tau$ on one of our smaller physical-volume ensembles. Eq. 36 requires as input the infinite-volume limit of $a_{\rm lat}/\ell$, which is needed to set $\ell_{V}/\ell$. Although all of our ensembles are consistent with the classical de Sitter solution, at least within a chosen fit window, this implies a particular finite-volume value of $a_{\rm lat}/\ell$. Because the lattice volumes at which we are simulating are rather small in Planck units, and the classical de Sitter solution that is used to fix $a_{\rm lat}/\ell$ only applies at large volumes, we must extrapolate our values of $a_{\rm lat}/\ell$ to the infinite-volume limit before we can use them as conversion factors. When we use the appropriate infinite-volume extrapolated values of $a_{\rm lat}/\ell$ in Eq.  36, the behavior of $\Lambda$ versus $\tau$ on a finite volume shows significant deviations from the pure de Sitter solution, where a constant $\Lambda$ is expected. We find that these results are broadly consistent across lattice spacings and that the Euclidean time dependence of the effective cosmological constant does not appear to be a lattice artifact. In the remainder of this work we show that these deviations can be described by nontrivial vacuum dynamics and that they pass a number of important consistency checks in this interpretation.
@@
In order to satisfy diffeomorphism invariance, the evolution must obey covariant energy conservation. The various models of Solà et al. [26, 29] differ in the way that this condition is satisfied. The lattice uniquely singles out one of these, as we show. We find that the vacuum pressure adjusts itself so that the vacuum sector (interpreted as an exotic form of matter) independently satisfies covariant energy conservation, $T^{\mu\nu}_{\ \ ;\,\mu}=0$, leading to
@@
Eq. (55) provides an alternative determination of $\nu$ within this model for running vacuum energy under the assumption that we can neglect the terms proportional to $\tilde{\nu}$. The agreement between the different determinations of $\nu$ provides a non-trivial check that the model is consistent with lattice data across observables.
@@ IV.2 Determining model parameters from the lattice
An alternative determination of $\nu^{\prime}$ allows for a powerful crosscheck of the model. Forming the combination
@@ IV.5 Alternative calculation of $\nu^{\prime}$
As discussed in subsection IV.2, the FLRW equations suggest an alternative determination of $\nu^{\prime}$ from the one presented in subsection IV.3. The second FLRW equation [Eq. (45)], when combined with the assumption of covariant energy conservation, leads to Eq. (59), which can be solved for $\nu^{\prime}$ to obtain
@@
A combined fit to the data of Fig. 22 to the form of Eq. (62) leads to a good fit, with a $\chi^{2}/{\rm d.o.f.}$ of $0.937$ corresponding to a $p$-value of $0.514$. The values of the fit parameters, $A^{\prime}=0.136(28)$ and $B^{\prime}=0.132(10)$, are in excellent agreement with those from the first method of determining $\nu^{\prime}$, as can be seen by comparing to the results of Table 7. This agreement between the two methods for determining $\nu^{\prime}$ is a strong test of the model defined by Eqs. (46) and (48). Even if one were inclined to dismiss a non-zero $\nu^{\prime}$ as a lattice artifact, the good agreement between both methods, ensuring covariant energy conservation, is a powerful check that the emergent behavior of the vacuum in EDT is consistent with one of the key properties expected of a theory of gravity.

### v2

@@ Abstract
We study the behavior of the vacuum in Euclidean dynamical triangulations (EDT). Algorithmic improvements and better lattice spacing determinations allow us to test the properties of the emergent de Sitter geometries of our simulations to higher precision than previously possible. Although the agreement with de Sitter is good, the improved precision reveals deviations that can be interpreted as non-trivial vacuum dynamics, well-described by a cosmological constant that runs with scale. The simulations show that the dominant running is quadratic and that the scale can be identified with the Hubble rate. Several key cross-checks support this picture, including consistent results across multiple lattice spacings and the fact that the null energy condition is not violated. The parameters of the running are fully determined by simulations, enabling predictions when extrapolated to the scales relevant for our universe. This leads to a model for dark energy that is compatible with current observations, but which predicts deviations from $\Lambda$CDM at the ${\cal O}(10^{-3})$ level in cosmological observables that could be tested with future improvements in precision measurements.
@@ I Introduction
The machinery introduced here and in Ref. [19] thus enables a strong test of our simulation results against expectations from the semiclassical de Sitter solution. This new precision allows us to show that despite the good overall agreement with this picture, there are deviations from semiclassical de Sitter that can be associated with nontrivial vacuum dynamics. This dynamics is well-described by a power-law running of the cosmological constant, where the renormalization scale is associated with the Hubble scale. This coincides with one of the simplest of a family of running vacuum models studied by Solà and collaborators [24, 25, 26, 27, 28, 29, 30]. These models have been compared to observations, thus constraining the free parameters of the models and testing whether the new physics implied by these models can resolve the emerging tensions in observational cosmology [28, 31]. We are able to turn this around by fitting lattice data directly to a specific running vacuum model, thus determining the free parameters of the model from first principles via our calculations. We find that only one parameter, a dimensionless coupling, is compatible with a nonzero value. The evidence for this parameter being different from zero is compelling, with high statistical significance and good agreement across four lattice spacings, making it unlikely that this effect is merely a lattice artifact.
The sign of the effect seen here ensures that it does not violate the null-energy condition. There are strong constraints on theories that are inconsistent with this condition [32, 33, 34], so the fact that our lattice result does not run afoul of it is a significant point in its favor. Although our results are found on lattices that are no larger than several Planck volumes, it is interesting to ask what would happen if the model singled out by this work was assumed to hold throughout cosmic history. One consequence is that the running of the vacuum energy would lead to a large vacuum energy density in the early universe compared to the current era. It turns out that the computed size of the coupling of the model ensures that the effects of the vacuum on the evolution of the universe remain subdominant throughout cosmic history, thus making it compatible with present observational bounds. Nonetheless, the corrections to standard cosmology would not be negligible, with the effects of the running vacuum giving rise to deviations from the $\Lambda$CDM model at the part-per-mil level across many cosmological observables, including the dark energy equation of state.
@@ II.1 Euclidean Dynamical Triangulations
$Z_{E}=\sum_{T}\frac{1}{C_{T}}\left[\prod_{j=1}^{N_{2}}{\cal O}(t_{j})^{\beta}\right]e^{-S_{ER}},$ (3)
where $N_{i}$ is the number of simplices of dimension $i$, and $C_{T}$ is a symmetry factor that divides out the number of equivalent ways of labeling the vertices in the triangulation $T$. The term in brackets in Eq. (19) is a nonuniform measure term [20, 35], where the product is over all triangles, and ${\cal O}(t_{j})$ is the order of triangle $j$, i.e. the number of four-simplices to which the triangle belongs. This corresponds in the continuum to a nonuniform weighting of the measure in Eq. (1) by $\prod_{x}\sqrt{g}^{\beta}$, where we leave $\beta$ an adjustable parameter in the simulations.
@@
$S_{E}=-\frac{\sqrt{3}}{2}\pi\kappa N_{2}+N_{4}\left(\kappa\frac{5\sqrt{3}}{2}\mbox{arccos}\frac{1}{4}+\frac{\sqrt{5}}{96}\lambda\right).$ (6)
We rewrite the Einstein-Regge action in the simple form
@@ II.3 The phase diagram and recovering physical results
The tuning of the bare lattice parameters to the phase transition is performed as follows: The emergent shape of the lattice geometries differs markedly between the two phases, allowing us to use a measurement of the shape to identify the transition. To do this, we introduce the shelling function $N_{4}^{\rm shell}(\tau)$, which is defined as the number of four-simplices lying exactly a geodesic distance $\tau$ away from a source four-simplex, comprising a spherical shell with the source at its center. This function is averaged over multiple random sources on a given configuration and over all configurations of an ensemble. We account for autocorrelation errors present in the data for the shelling function by blocking the data before averaging. We study the variation of the error with block size, increasing the block size until the standard error obtained from a jackknife average over configurations no longer increases. The height of the peak (as a function of $\tau$) of the shelling function $N_{4,\mathrm{peak}}^{\rm shell}$ varies significantly across the phase transition and thus serves as a good order parameter for studying the phase diagram.
@@
As described in Ref. [19], there is a “knee” in the plots at each volume, just before the slope becomes large and negative, that marks the onset of the phase transition. We take a fixed value of $\kappa_{2}$ to define our nominal lattice spacing, and we then tune $\beta$ by matching the rescaled peak heights in the region just to the left of the phase transition. The $\beta$ value is chosen so that the tuned point is sufficiently far from the phase transition that there is no evidence of the tunneling between metastable states that is expected in the vicinity of a first-order phase transition. Adjusting the rescaling of the peak height so that the location of the “knee” is at the same height across different volumes leads to a $D_{H}$ close to the physical value of four. We thus assume that the peak height rescaling is four-dimensional, and we choose $\beta$ as a function of volume so that the rescaled peak heights match the rescaled peak height of one of our smaller volumes near the transition. The tuned value is chosen so that a broad range of volumes can be adjusted so as to match the peak height after the rescaling. If the volume is chosen too small at a given lattice spacing, it is not possible to reach the tuned peak height, and indeed, for too small of a physical volume, the identification with semiclassical geometry with $D_{H}=4$ is not possible. The choice of the tuned peak height varies smoothly across the different lattice spacings, with its value given by the cyan band in each plot of Fig. 2. It is the 8k volume at $\kappa_{2}=2.45$, the 12k volume at $\kappa_{2}=3.0$, the 16k volume at $\kappa_{2}=3.4$, and the 32k volume at $\kappa_{2}=3.8$. As we show in Section III, this collection of ensembles passes a broad set of consistency checks, including the recovery of the semiclassical limit, suggesting that this tuning procedure is a valid one.
Figure 2: Rescaled peak heights for multiple volumes at four lattice spacings. The horizontal cyan line represents the peak height to which the tuning is performed for all volumes at that lattice spacing. The volume picked for each lattice spacing is the 8k volume at $\kappa_{2}=2.45$, the 12k volume at $\kappa_{2}=3.0$, the 16k volume at $\kappa_{2}=3.4$, and the 32k volume at $\kappa_{2}=3.8$.
@@ III.1 The classical de Sitter solution
where $\ell$ is the dual lattice spacing and $\ell_{V}$ is the unit length derived from taking the cube root of the normalization of the shelling function. These units of measurement can be related via a comparison of the total four-volume of a thin slice of de Sitter space measured either by summing the volumes of four-simplices making up that slice, or by taking the integral over a four-dimensional shell. This leads to the relation
@@ III.3.1 $a_{\rm lat}/\ell$ from Method 1
Figure 4: The rescaled shelling function $n_{4}(\rho)$ at different lattice spacings. The black curve is the classical de Sitter solution. All curves have been rescaled to overlap in the region that matches the classical solution. The asymmetry at large Euclidean time decreases as the lattice spacing gets finer, so that the lattice results approach the classical curve. The $\beta=1.5$ ensemble shown here has a volume of 4k, and all the others have a volume of 32k.
@@ III.4 Determining Newton’s constant and setting the lattice spacing
Table 17, Table 18, Table 19, and Table 20 show the results for $|s_{G}|$ from all of our scaling studies. The first error is statistical, and the second is an estimate of the systematic error due to varying the $\kappa_{4}$ data points included in the fits. Adding these errors in quadrature and constructing $G/a_{\rm lat}^{2}$ from Eq. (28) leads to the values of $G/a_{\rm lat}^{2}$ shown in Fig. 11, plotted as a function of $1/N_{4}$. We have determined $G/a_{\rm lat}^{2}$ at multiple volumes at each lattice spacing, allowing us to use the infinite-volume limit of $G/a_{\rm lat}^{2}$, evaluated separately for each $\kappa_{2}$ value, to determine a relative lattice spacing. In previous work [21] there was only one lattice spacing available with multiple volumes, so this was not possible. In that case, the relative lattice spacing was determined using the diffusion process of a random walk. In revisiting the return probability to obtain the relative lattice spacing on our new ensembles, we find that there is significant ambiguity in rescaling the return probability curve, given how much the shape changes with lattice spacing. Using our $G/a_{\rm lat}^{2}$ for this purpose removes that ambiguity, while also being systematically improvable with more computing resources. Compared to [21], the relative lattice spacings change by up to a factor of $\sim 1.7$. We take the $\kappa_{2}=3.0$ ensembles to define our fiducial lattice spacing, so that $a_{\rm rel}$ at this lattice spacing is by definition one. The other lattice spacings are measured relative to this, being inversely proportional to the square root of the ratio of their $G/a_{\rm lat}^{2}$ values. The value of $G/a_{\rm lat}^{2}$ at $\kappa_{2}=3.0$ then sets our absolute lattice spacing, allowing us to convert our measurements to units of the Planck length.
@@ IV.1 Theory and motivation
Our lattice geometries are well-described by the semi-classical Euclidean de Sitter solution, which is solid evidence that EDT could define a non-perturbative formulation of quantum gravity. Interestingly, a more careful study shows that there are corrections to this picture. As pointed out in Section III.3.2, we find that the cosmological constant is not strictly constant as a function of Euclidean time, particularly on our smaller volume ensembles. A deviation from the expected classical scaling of the geometries with volume could provide a window to quantum gravitational effects, assuming that they survive a continuum limit. Although such deviations might be treated as finite-size effects to be extrapolated away in a flat space lattice QCD calculation, in quantum gravity the dynamics of the universe at small sizes is an interesting question in its own right. Figure 12 shows the cosmological constant inferred from Eq. 36 as a function of Euclidean time $\tau$ on one of our smaller physical-volume ensembles. Eq. 36 requires as input the infinite-volume limit of $a_{\rm lat}/\ell$, which is needed to set $\ell_{V}/\ell$. Although all of our ensembles are consistent with the classical de Sitter solution, at least within a chosen fit window, this implies a particular finite-volume value of $a_{\rm lat}/\ell$. Because the lattice volumes at which we are simulating are rather small in Planck units, and the classical de Sitter solution that is used to fix $a_{\rm lat}/\ell$ only applies at large volumes, we must extrapolate our values of $a_{\rm lat}/\ell$ to the infinite-volume limit before we can use them as conversion factors. When we use the appropriate infinite-volume extrapolated values of $a_{\rm lat}/\ell$ in Eq.  36, the behavior of $\Lambda$ versus $\tau$ on a finite volume shows significant deviations from the pure de Sitter solution, where a constant $\Lambda$ is expected. We find that these results are broadly consistent across lattice spacings and that the Euclidean time dependence of the effective cosmological constant does not appear to be a lattice artifact. In the remainder of this work we show that these deviations can be described by nontrivial vacuum dynamics and that they pass a number of important consistency checks in this interpretation.
@@
In order to satisfy diffeomorphism invariance, the evolution must obey covariant energy conservation. The various models of Solà et al. [26, 31] differ in the way that this condition is satisfied, postulating different ways that the vacuum energy might transfer energy to the matter sector. In our pure gravity simulations, we do not see evidence for an emergent matter sector in the cosmological dynamics, so that the vacuum sector independently satisfies covariant energy conservation, $T^{\mu\nu}_{\ \ ;\,\mu}=0$, leading to
@@
Eq. (55) provides an alternative determination of $\nu$ within this model for running vacuum energy under the assumption that we can neglect the terms proportional to $\tilde{\nu}$.
@@ IV.2 Determining model parameters from the lattice
An alternative determination of $\nu^{\prime}$ allows for a useful crosscheck of the model. Forming the combination
@@ IV.5 Alternative calculation of $\nu^{\prime}$
As discussed in subsection IV.2, the FLRW equations suggest an alternative determination of $\nu^{\prime}$ from the one presented in subsection IV.3. The second FLRW equation [Eq. (45)]can be solved for $\nu^{\prime}$ to obtain
@@
A combined fit to the data of Fig. 22 to the form of Eq. (62) leads to a good fit, with a $\chi^{2}/{\rm d.o.f.}$ of $0.937$ corresponding to a $p$-value of $0.514$. The values of the fit parameters, $A^{\prime}=0.136(28)$ and $B^{\prime}=0.132(10)$, are in excellent agreement with those from the first method of determining $\nu^{\prime}$, as can be seen by comparing to the results of Table 7. This agreement between the two methods for determining $\nu^{\prime}$ is a useful test of the model defined by Eqs. (46) and (48).